41. The budget process involves doing all of the following except                    

              A.    executing plans to achieve the goals                                          

              B.    firing all managers who fail to achieve operational goals specified in the budget       

              C.    periodically comparing actual results with the goals                                 

              D.    establishing specific goals                                     

42. The process of developing budget estimates by requiring all levels of management to estimate sales, production, and other operating data as though operations were being initiated for the first time is referred to as

              A.    continuous budgeting                                          

              B.    master budgeting                                                

              C.    zero-based budgeting                                                

              D.    flexible budgeting                                                

43. Jase Manufacturing Co.'s static budget at 10,000 units of production includes $40,000 for direct labor and $4,000 for electric power. Total fixed costs are $24,000. At 12,000 units of production, a flexible budget would show

              A.    variable costs of $44,000 and $24,000 of fixed costs                                 

              B.    variable and fixed costs totaling $68,000                             

              C.    variable costs of $52,800 and $24,000 of fixed costs                                 

              D.    variable costs of $52,800 and $29,000 of fixed costs                                 

44. If budgeted beginning inventory is $8,000, budgeted ending inventory is $9,400, and budgeted cost of goods sold is $10,260, budgeted production should be

              A.    $9,600                                                   

              B.    $11,550                                                 

              C.    $11,660                                                 

              D.    $1,400                                                   

45. Stephanie Corporation sells a single product. Budgeted sales for the year are anticipated to be 697,000 units, estimated beginning inventory is 106,000 units, and desired ending inventory is 88,000 units. The quantities of direct materials expected to be used for each unit of finished product are given below.

                     Material A: 0.50 lb. per unit @ $0.60 per pound                                 

                     Material B: 1.00 lb. per unit @ $1.53 per pound                                 

                     Material C: 1.20 lb. per unit @ $1.18 per pound                                 

                     The dollar amount of material A used in production during the year is              

              A.$203,700 B.$1,038,870 C.$209,100    D.$961,464
Answer #1
41
The budget process involves doing all of the following except firing all managers who fail to achieve operational goals specified in the budget
Option B is correct
42
The process of developing budget estimates as if were being initiated for the first time is referred to as zero-based budgeting
Option C is correct
43
Variable costs = (40000+4000)/10000*12000= $52,800
Fixed costs = $24,000
Variable costs of $52,800 and $24,000 of fixed costs      
Option C is correct
44
Budgeted production = Budgeted cost of goods sold+Budgeted ending inventory-Budgeted beginning inventory
Budgeted production = 10260+9400-8000= $11,660
Option C is correct
45
Budgeted production = 697000+88000-106000= 679000
Dollar amount of material A = 679000*0.50*0.60= $203,700
Option A is correct